# Unveiling the jet angular broadening with photon-tagged jets in high-energy nuclear collisions

### v2

@@ III Results and discussions
where $p_{T1}$ and $p_{T2}$ represent the transverse momentum of the leading and subleading subjets, respectively, while $R$ is the jet size. $\eta$ and $\phi$ are the pseudorapidity and azimuthal angle of the two subjets. We set $z_{\text{cut}}=0.2$ and $\beta=0$ in line with the CMS measurement CMS:2024zjn . The theoretical results for the modification of the groomed radius of $\gamma$+jets show a moderate enhancement at $R_{g}>0.15$, consistent with the CMS data, which may indicate jet angular broadening. In contrast, we observe that the medium modification pattern for inclusive jets shows a notable narrowing, diverging from the behavior of $\gamma$+jets. Additionally, by utilizing Jet-by-Jet matching as shown in the lower panel of Fig. 7, we can see an enhancement of the PbPb/pp ratio at $R_{g}>0.1$ for both inclusive jets and $\gamma$+jets. This observation is analogous to what we discussed earlier regarding jet girth in Fig. 3. Both inclusive jets and $\gamma$+jets become broader when they pass through the QGP. However, the selection bias reduces the broadening effect seen in the $R_{g}$ distribution of inclusive jets and $\gamma$+jets in nucleus-nucleus collisions. Despite this suppression, $\gamma$+jets exhibit a slight broadening effect, which is less influenced by selection bias compared to inclusive jets.
@@
Figure 8: (Color online) Medium modification of the angle between jet axes of $\gamma$+jet and inclusive jet in $0-10\%$ PbPb collisions relative to pp at $\sqrt{s_{NN}}=5.02$ TeV. The results are compared with the measurements of the inclusive jet by the ALICE ALICE:2023dwg . We also plot the results using the Jet-by-Jet matching (JBJ) as a comparison in the lower panel.
@@
In Fig. 8, we compare the medium modification of the angle between jet axes, denoted as $\Delta R_{\rm axis}$, for $\gamma$+jets and inclusive jets in $0-10\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV. The PbPb/pp represents the ratio of normalized $\Delta R_{\rm axis}$ distributions in PbPb to pp collisions. The $\Delta R_{\rm axis}$ quantifies the angular distance between the directions of jet axes reconstructed using both the standard and Winner-Take-All (WTA) recombination schemes Cal:2019gxa :
$\Delta R_{\rm axis}\equiv\sqrt{(y_{\rm axis}^{\rm WTA}-y_{\rm axis}^{\rm Std})^{2}+(\phi_{\rm axis}^{\rm WTA}-\phi_{\rm axis}^{\rm Std})^{2}}.$ (11)
@@
Here, $y_{\rm axis}$ and $\phi_{\rm axis}$ represent the rapidity and azimuthal angle of the jet axes, respectively. The standard jet axis is derived by summing the momenta of all constituent particles of the anti-$k_{T}$ jet using a clustering algorithm known as the E-scheme. In contrast, the WTA axis is obtained by reclustering the anti-$k_{T}$ jets with the WTA recombination scheme Bertolini:2013iqa . This WTA axis is generally aligned with the most energetic constituent within the jets and minimizes the impact of soft particles on the direction of the jet axis. Therefore, the distance between the standard and WTA axes can be used to investigate the role that soft particles play in the evolution of jets in a medium. Recent studies have measured the medium modification of the inclusive jet $\Delta R_{\rm axis}$ in heavy-ion collisions, as reported by the ALICE collaboration ALICE:2023dwg . Experimental results indicate that inclusive jets show narrower $\Delta R_{\rm axis}$ distributions within the range $6040<p_{T}^{\rm jet}<$ 60 GeV in PbPb collisions compared to pp collisions. Default calculations of the $\Delta R_{\rm axis}$ modification for inclusive jets are consistent with the ALICE data. We also assess the nuclear modification of the $\gamma$+jet $\Delta R_{\rm axis}$ in the same collision system, applying different kinematic cuts ($x_{j\gamma}>0.4$, $p_{T}^{\gamma}>100$ GeV). Our calculations reveal a slight enhancement in the PbPb/pp ratio of the $\gamma$+jet $\Delta R_{\rm axis}$ distribution in PbPb collisions for $\Delta R_{\rm axis}>0.06$ compared to pp collisions. This suggests that angular jet broadening within the QGP can also be measured through the $\Delta R_{\rm axis}$ distribution of $\gamma$+jets in heavy-ion collisions. In the lower panel of Fig. 8, the Jet-by-Jet matching results exhibit clear broadening without the influence of selection bias both for $\gamma$+jet and inclusive jet, which indicates that selection bias suppresses the broadening effect of $\Delta R_{\rm axis}$ distribution in nucleus-nucleus collisions. Especially, the influence of selection bias reverses broadening to narrowing in the $\Delta R_{\rm axis}$ medium modification pattern for inclusive jets. Therefore, the differing nuclear modification patterns of $R_{g}$ and $\Delta R_{\rm axis}$ for inclusive jets and $\gamma$+jets could be of significant interest for future experimental verification. The findings presented in this paper will aid in interpreting recent measurements at the LHC and will be beneficial for future studies focusing on the intrinsic modification of jet substructure in heavy-ion collisions.
@@ IV Summary
In summary, we present a theoretical study on the angular structure of $\gamma$+jets in high-energy nuclear collisions at the LHC. We utilize the PYTHIA8 to provide the initial production of $\gamma$+jets and employ a transport approach to simulate the in-medium jet energy loss in nucleus-nucleus collisions. We carry out the medium modification of $\gamma$+jet girth in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=$ 5.02 TeV, which shows a satisfactory agreement with the recently reported CMS data.
Furthermore, we investigate the influence of selection bias when choosing different $x_{j\gamma}$ cuts. With the help of the Jet-by-Jet matching method, we explore the connection between selection bias and kinematic requirements in event selection. Importantly, we quantitatively demonstrate that $\gamma$+jets will provide significant advantages to reduce the selection bias and can effectively collect jets sufficiently quenched in PbPb collisions compared to the inclusive jets. We also discuss the contributions of medium-induced gluon radiation and the medium response to the broadening of jet angular substructure in PbPb collisions. Additionally, to make the discussion more general, we also study the modification patterns of $R_{g}$ and $\Delta R_{\rm axis}$ for inclusive jets and $\gamma$+jets in PbPb collisions, which show slight broadening for $\gamma$+jets but evident narrowing for inclusive jets.
This theoretical study provides new insights into the extensive measurements focusing on intra-jet broadening in heavy-ion collisions ALargeIonColliderExperiment:2021mqf ; ATLAS:2022vii ; ALICE:2018dxf ; ALICE:2023dwg ; Ehlers:2022dfp ; ATLAS:2023hso , as well as the acoplanarity broadening recently observed for lower $p_{T}$ jets STAR:2023pal ; STAR:2023ksv ; ALICE:2023qve ; ALICE:2023jye . We anticipate further measurements of $\gamma$+jet substructure in heavy-ion collisions, which may provide critical constraints on current theoretical frameworks regarding jet-medium interactions.
@@ Acknowledgements.
This research is supported by the Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008, and the National Natural Science Foundation of China with Project Nos. 11935007 and 12035007. S. W. is supported by the Open Foundation of Key Laboratory of Quark and Lepton Physics (MOE) No. QLPL2023P01 and the Talent Scientific Star-up Foundation of China Three Gorges University (CTGU) No. 2024RCKJ013.

### v4

@@ III Results and discussions
where $p_{T1}$ and $p_{T2}$ represent the transverse momentum of the leading and subleading subjets, respectively, while $R$ is the jet size. $\eta$ and $\phi$ are the pseudorapidity and azimuthal angle of the two subjets. We set $z_{\text{cut}}=0.2$ and $\beta=0$ in line with the CMS measurement Hayrapetyan et al. (2025b). The theoretical results for the modification of the groomed radius of $\gamma$+jets show a moderate enhancement at $R_{g}>0.15$, consistent with CMS data, which may indicate jet angular broadening. In contrast, we observe that the medium modification pattern for inclusive jets shows a notable narrowing, diverging from the behavior of $\gamma$+jets. Additionally, by utilizing Jet-by-Jet matching as shown in the lower panel of Fig. 7, we can see an enhancement of the PbPb/pp ratio at $R_{g}>0.1$ for both inclusive jets and $\gamma$+jets. This observation is analogous to what we discussed earlier regarding jet girth in Fig. 3. Both inclusive jets and $\gamma$+jets become broader when they pass through QGP. However, selection bias reduces the broadening effect seen in the $R_{g}$ distribution of inclusive jets and $\gamma$+jets in nucleus-nucleus collisions. Despite this suppression, $\gamma$+jets exhibit a slight broadening effect, which is less influenced by selection bias compared to inclusive jets.
@@
Figure 8: (Color online) Medium modification of the angle between jet axes of $\gamma$+jet and inclusive jet in $0-10\%$ PbPb collisions relative to pp at $\sqrt{s_{NN}}=5.02$ TeV. The results are compared with the measurements of the inclusive jet by the ALICE Acharya et al. (2023). We also plot the results using the Jet-by-Jet matching (JBJ) as a comparison in the lower panel.
@@
In Fig. 8, we compare the medium modification of the angle between jet axes, denoted as $\Delta R_{\rm axis}$, for $\gamma$+jets and inclusive jets in $0-10\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV. The PbPb/pp represents the ratio of normalized $\Delta R_{\rm axis}$ distributions in PbPb to pp collisions. The $\Delta R_{\rm axis}$ quantifies the angular distance between the directions of jet axes reconstructed using both the standard and Winner-Take-All (WTA) recombination schemes Cal et al. (2020):
$\Delta R_{\rm axis}\equiv\sqrt{(y_{\rm axis}^{\rm WTA}-y_{\rm axis}^{\rm Std})^{2}+(\phi_{\rm axis}^{\rm WTA}-\phi_{\rm axis}^{\rm Std})^{2}}$ (10)
@@
Here, $y_{\rm axis}$ and $\phi_{\rm axis}$ represent the rapidity and azimuthal angle of the jet axes, respectively. The standard jet axis is derived by summing the momenta of all constituent particles of the anti-$k_{T}$ jet using a clustering algorithm known as the E-scheme. In contrast, the WTA axis is obtained by reclustering the anti-$k_{T}$ jets with the WTA recombination scheme Bertolini et al. (2014). This WTA axis is generally aligned with the most energetic constituent within the jets and minimizes the impact of soft particles on the direction of the jet axis. Therefore, the distance between the standard and WTA axes can be used to investigate the role that soft particles play in the evolution of jets in a medium. Recent studies have measured the medium modification of the inclusive jet $\Delta R_{\rm axis}$ in heavy-ion collisions, as reported by the ALICE collaboration Acharya et al. (2023). Experimental results indicate that inclusive jets show narrower $\Delta R_{\rm axis}$ distributions within the range $40<p_{T}^{\rm jet}<60$ GeV in PbPb collisions compared to pp collisions. Default calculations of the $\Delta R_{\rm axis}$ modification for inclusive jets are consistent with the ALICE data. We also assess the nuclear modification of the $\gamma$+jet $\Delta R_{\rm axis}$ in the same collision system, applying different kinematic cuts ($x_{j\gamma}>0.4$, $p_{T}^{\gamma}>100$ GeV). Our calculations reveal a slight enhancement in the PbPb/pp ratio of the $\gamma$+jet $\Delta R_{\rm axis}$ distribution in PbPb collisions for $\Delta R_{\rm axis}>0.06$ compared to pp collisions. It suggests that angular jet broadening within QGP can also be measured through the $\Delta R_{\rm axis}$ distribution of $\gamma$+jets in heavy-ion collisions. In the lower panel of Fig. 8, the Jet-by-Jet matching results exhibit evident broadening without the influence of selection bias, both for $\gamma$+jet and inclusive jet, which indicates that selection bias suppresses the broadening effect of $\Delta R_{\rm axis}$ distribution in nucleus-nucleus collisions. Especially, the influence of selection bias reverses broadening to narrowing in the $\Delta R_{\rm axis}$ medium modification pattern for inclusive jets. Therefore, the differing nuclear modification patterns of $R_{g}$ and $\Delta R_{\rm axis}$ for inclusive jets and $\gamma$+jets could be of significant interest for future experimental verification. The findings presented in this paper will aid in interpreting recent measurements at LHC and will be beneficial for future studies focusing on the intrinsic modification of jet substructure in heavy-ion collisions.
@@ IV Summary
In summary, we present a theoretical study on the angular structure of $\gamma$+jets in high-energy nuclear collisions at the LHC. We utilize PYTHIA8 to provide the initial production of $\gamma$+jets and employ a transport approach to simulate the in-medium jet energy loss in nucleus-nucleus collisions. We carry out the medium modification of $\gamma$+jet girth in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV, which shows a good agreement with the recently reported CMS data.
Furthermore, we investigate the influence of selection bias when choosing different $x_{j\gamma}$ cuts. With the help of the Jet-by-Jet matching method, we explore the connection between selection bias and kinematic requirements in event selection. Importantly, we quantitatively demonstrate that $\gamma$+jets will provide significant advantages to reduce selection bias and can effectively collect jets sufficiently quenched in PbPb collisions compared to the inclusive jets. We also discuss the contributions of medium-induced gluon radiation and medium response to the broadening of jet angular substructure in PbPb collisions. Additionally, to make the discussion more general, we also study the modification patterns of $R_{g}$ and $\Delta R_{\rm axis}$ for inclusive jets and $\gamma$+jets in PbPb collisions, which show slight broadening for $\gamma$+jets but significant narrowing for inclusive jets.
This theoretical study provides new insights into the extensive measurements focusing on intra-jet broadening in heavy-ion collisions Acharya et al. (2022); Aad et al. (2023a); Acharya et al. (2018, 2023); Ehlers (2022); Aad et al. (2023b), as well as the acoplanarity broadening recently observed for lower $p_{T}$ jets Aboona et al. (2025a, b); Acharya et al. (2024a, b). We anticipate further measurements of $\gamma$+jet substructure in heavy-ion collisions, which may provide critical constraints on current theoretical frameworks regarding jet-medium interactions.
@@ Acknowledgements.
This research is supported by the National Natural Science Foundation of China with Project Nos. 12535010. S.W. is supported by the Open Foundation of Key Laboratory of Quark and Lepton Physics (MOE) No. QLPL2023P01 and the Talent Scientific Star-up Foundation of China Three Gorges University (CTGU) No. 2024RCKJ013.